OpenStudy (anonymous):

There are 40 people at a reception for international students. The students, boys and girls, are from Peru and China. 18 of all the students are boys, and 12 of all the students are Peruvian. 17 of the Chinese are girls. Given that a student is a boy, what is the probability that he is Chinese?

OpenStudy (amistre64):

P(AnB) = P(A) * P(B)

OpenStudy (anonymous):

i don't understand?

OpenStudy (amistre64):

i might have read that wrong ... there is a given in it :)

OpenStudy (amistre64):

There are 18 of all the students are boys, and 12 of all the students are Peruvian. 17 of the Chinese are girls. 40 people total 18 out of 40 are boys P(China) = P(not Peru) = (40-12)/40 = 28/40 28 -17 ----- 11 that leaves 11 chinese that are boys; out of 18 boys 11/18 is my best guess

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

yep